# Fast-weight Product Key Memory

### v1

@@ 3 Fast-weight Product Key Memory
The standard PKM is designed as a slow-weight channel mixer. In other words, its parameters are updated via a global objective (next-token prediction) over the entire training corpus but remain frozen during inference. Consequently, while it can store general semantic knowledge, it cannot adapt to new inputs or remember immediate contexts.
We propose to transform PKM into Fast-weight PKM (FwPKM). We redesign the module to update its parameters dynamically at both training and inference time, by optimizing a local objective. This enables FwPKM to function as a high-fidelity episodic memory that captures key-value associations from the input stream itself.
3.1 Fast Weights
In standard neural networks, knowledge is stored in “slow weights” – parameters $\phi$ that are optimized over a massive training dataset but frozen after training. While effective for storing general knowledge in the dataset, slow weights lack the ability to rapidly adapt to new contexts. The concept of “fast weights” (FW-Hinton; FW-Schmidhuber) addresses this problem by introducing a set of parameters that change dynamically according to every new input. The fast-weight parameters $\theta$ can be thought of as the storage of episodic memory.
Recent works like Test-Time Training (TTT, TTT) demonstrate that a fast-weight module can be implemented as a neural model $f(\cdot;\theta)$ and its parameters are updated by minimizing an MSE objective of an input sequence $\mathbf{h}_{1},\cdots,\mathbf{h}_{T}$. Formally, for an input $\mathbf{h}_{t}$ in the sequence, the model performs a gradient descent step to minimize the MSE of reconstructing $\mathbf{h}_{t}$ from its corrupted version $\tilde{\mathbf{h}}_{t}$ (e.g. via a low-rank projection):
$\theta^{\prime}$ $=\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{MSE}}(f(\tilde{\mathbf{h}}_{t};\theta),\mathbf{h}_{t})$ (10)
$=\theta-\eta\nabla_{\theta}\|\mathbf{h}_{t}-f(\tilde{\mathbf{h}}_{t};\theta)\|^{2}_{2},$ (11)
where $\eta$ is the learning rate. Through this optimization process, fast weights $\theta$ learn to encode information of the input sequence, at both training and inference time.
Following TTT, we implement fast weights by performing gradient descent updates on a local MSE objective during the forward pass. Specifically, we treat the key and value matrices as fast-weight parameters $\theta=\{K,V\}$. These parameters are updated via chunk-level gradient descent to minimize the reconstruction error of a chunk of query-value pairs.
Formally, for a chunk of size $C$, we generate query and target value inputs $\{(\mathbf{q}_{t},\mathbf{v}_{t})\}_{t=1}^{C}$ using slow-weight projections of the hidden states [^1] We abuse the subscript of a tensor to index both the sequence dimension and the feature dimension(s). We use $i,j$ to index the feature dimensions and $t$ to index the sequence dimension..
$\mathbf{q}_{t}$ $=\mathtt{Linear}^{q}_{\phi}(\mathtt{RMSNorm}^{q}_{\phi}(\mathbf{h}_{t}))$ (12)
$\mathbf{v}_{t}$ $=\mathtt{Linear}^{v}_{\phi}(\mathtt{RMSNorm}^{v}_{\phi}(\mathbf{h}_{t})).$ (13)
The model then computes a prediction $\hat{\mathbf{v}}_{t}$ using the current fast weights $\theta$ and updates it to minimize the Mean Square Error (MSE) between the prediction and target. We set the learning rate $\eta$ to 1.0 and multiply the loss with a constant factor of $0.5$.
$\hat{\mathbf{v}}_{t}$ $=\mathtt{PKM}(\mathbf{q}_{t};\theta)$ (14)
$\theta^{\prime}$ $=\theta-\textstyle\sum_{t=1}^{C}\nabla_{\theta}\frac{1}{2}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}}_{t},\mathbf{v}_{t}).$ (15)
3.2 Memorization Optimization
3.2.1 MSE loss
The optimization goal is to “rewrite” the memory such that retrieving with query $\mathbf{q}_{t}$ yields the target value $\mathbf{v}_{t}$. We employ the Mean Squared Error (MSE) as our objective due to its favorable gradient properties for explicit memory rewriting.
The loss for a single sample is $\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}},\mathbf{v})=\frac{1}{2}\|\mathbf{v}-\hat{\mathbf{v}}\|^{2}_{2}$. Notably, we use a constant factor $0.5$ and a learning rate of $\eta=1.0$ as in Eq. 15. This is not arbitrary because it achieves the effect of “one-step rewriting”. The gradient w.r.t. the prediction $\hat{\mathbf{v}}$ is $(𝐯𝐯^)\nabla_{\hat{\mathbf{v}}}\frac{1}{2}\mathcal{L}_{\text{MSE}}$($\hat$$\mathbf$$\mathbf{v}$)=-($\mathbf$$\hat{\mathbf{v}}$). A single gradient step therefore updates the prediction directly to the target.
$\hat{\mathbf{v}}^{\prime}=\hat{\mathbf{v}}-1.0\cdot(-(\mathbf{v}-\hat{\mathbf{v}}))=\mathbf{v}.$ (16)
This allows FwPKM to instantly memorize the new key-value association.
In practice, we optimize the value matrix that produced the prediction, and the gradient w.r.t. the $i$-th value matrix row is instead given by
$\nabla_{V_{i}}\frac{1}{2}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}},\mathbf{v})$ $=-(\mathbf{v}-\hat{\mathbf{v}})s^{\prime}_{i}.$ (17)
3.2.2 Loss aggregation and gradient shaping
The last section showed the gradients for a single sample, but we process in chunks. As a result, a) an MSE loss will be calculated on every feature for every predicted value $\hat{\mathbf{v}}$ and b) multiple tokens may attempt to write the same value row simultaneously. In the following, we show the proper way to reduce the MSE losses in a chunk into a scalar loss for backpropagation and how to aggregate the gradients for a row in the value matrix.
Reducing MSE losses to a scalar by summation
Instead of using averaging reduction, we reduce the MSE losses of a chunk of samples by summing over the sample and feature dimensions. Otherwise the gradient w.r.t. a value matrix element $V_{i,j}$ will be proportional to the actual update signal with a constant of $1/(\texttt{num\_samples}\times D_{V})$.
Weighting gradient by row contribution
Let $N^{\text{read}}_{i}$ be the “row contribution”, i.e. the number of times value row $V_{i}$ is accessed within the current chunk. We scale the gradient for row $i$ by $1/N^{\text{read}}_{i}$.
$\nabla^{\text{agg}}_{V_{i}}=\frac{1}{N^{\text{read}}_{i}}\textstyle\sum_{t=1}^{C}\nabla_{V_{i}}\frac{1}{2}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}}_{t},\mathbf{v}_{t}).$ (18)
This averaging strategy acts as a consensus mechanism, ensuring that competitive memory writes from different tokens in the same chunk are balanced rather than accumulative.
Weighting MSE loss by token importance
While the averaging strategy mitigates excessive memory writing, it treats all writings to the same slot equally. We weigh each MSE loss by a “gating value” $g_{t}$ (defined in Section 3.4.4) to prioritize important updates within competitive writings. This scalar value represents the strength of FwPKM’s impact on the language model and the usefulness of its corresponding token.
No gradient clipping
Unlike standard training, we explicitly avoid gradient clipping to fast-weight updates. Since the target values $\mathbf{v}$ are unbounded, unclipped gradients are helpful for the memory to fully adapt to the scale of the target values.
3.3 Addressing Optimization
3.3.1 Marginal entropy loss
Sparse memory suffers from “memory collapsing” where the model learns to utilize only a small number of memory slots. PKM showed that a vanilla PKM of $512\times 512=262k$ slots can only reach $64.4\%$ slot usage when selecting effectively Top-$128$ slots per token, with a multi-head mechanism where $4$ heads $\times$ Top-$32$ per head =128=128 slots. The authors proposed applying batch normalization to queries and it raises the usage to $97.9\%$. However, we found the query normalization technique to become inefficient when the effective Top-$k$ is small (e.g. $1$ head with Top-$8$ slots per head), and a small effective Top-$k$ is shown to be important for building a performant FwPKM (See ablation study in Appendix B).
We counteract FwPKM’s memory collapsing by optimizing an auxiliary addressing objective based on marginal entropy maximization. The goal is to encourage the model to access all memory slots uniformly on average across a chunk, without forcing a uniform distribution for any single individual query. Formally, for each of FwPKM’s sub key set, let $\mathbf{s}^{\prime}_{t}\in\mathbb{R}^{\sqrt{N}}$ denote the normalized query-key scores after Top-$k$ selection for token $t$ in a chunk, where unselected indidces have score 0. We compute the marginal distribution $\bar{\mathbf{p}}$ representing the average slot usage over the chunk and define the addressing loss as the marginal entropy of $\bar{\mathbf{p}}\in\mathbb{R}^{\sqrt{N}}$:
$\bar{\mathbf{p}}$ $=\frac{1}{C}\textstyle\sum_{t=1}^{C}\mathbf{s}^{\prime}_{t}$ (19)
$\mathcal{L}_{\text{addr}}$ $=-H(\bar{\mathbf{p}})=\textstyle-\sum_{i=1}^{\sqrt{N}}\bar{p}_{i}\log\bar{p}_{i}.$ (20)
While minimizing the MSE loss optimizes FwPKM’s value matrix $V$ to store input values, minimizing the marginal entropy loss trains the key matrix $K$ to adapt to input queries’ distribution, making the key vectors cover the query representation space more effectively and uniformly.
3.3.2 IDW score
A query-key score $s_{i}$ in PKM is the dot product between a query and a key row $\mathbf{q}^{\top}K_{i}$. However, a key row can learn to change its magnitude to produce a larger score for a target query, without having to be “close” to the query in the representation space. Inverse distance weight (IDW, IDW) score is an alternative to dot-product score that produces a different key layout.
$s^{\text{IDW}}_{i}=-\log(\epsilon+\|\mathbf{q}-K_{i}\|_{2}^{2}),$ (21)
where $\epsilon=10^{-3}$ following the paper. Due to the use of Euclidean distance, gradients produced by IDW scores will push keys to be “prototypes” – centroids of query clusters. We found that IDW score yields better performance than dot-product score.
3.4 Target Value Construction
3.4.1 Value residual
Target values are produced by a slow-weight network $\mathbf{v}_{t}=\mathtt{Linear}^{v}_{\phi}(\mathtt{RMSNorm}^{v}_{\phi}(\mathbf{h}_{t}))$. These slow-weight parameters, however, only directly participate in the computation of the fast-weight MSE loss whose gradients cannot reach slow weights. We add a residual connection from the output of the value projection layer to the output of FwPKM, so the value projection parameters are on the forward/backward path of the LM loss, which provides learning signals to make them produce target values that are useful for next token prediction.
3.4.2 Lookahead value
In a similar spirit to using short convolution in linear attention variants (RWKV; GLA; Mamba), we pair queries with lookahead values when applying chunk-level updates. Essentially, we slightly modify the update rule in Eq. 15 by changing the timestep subscript of target value from $t$ to $t+1$. In this way, FwPKM associates each token’s key to its next token’s value, which provides more useful information for next token prediction.
$\hat{\mathbf{v}}_{t+1}$ $=\mathtt{PKM}(\mathbf{q}_{t};\theta)$ (22)
3.4.3 Target value normalization
We also found it useful to z-score normalize target values on the feature dimension. Although the absence of grad clipping already ensures that value matrix $V$ can adapt inputs of arbitrary scale, constraining the target values to have a mean of 0 and a standard deviation of $1$ improves training stability.
3.4.4 Gating
Next token prediction does not always rely on episodic memory, so we devise a gating mechanism to give the model the freedom of determining how much information is extracted from FwPKM outputs. Similar to the computation of query and value vectors, we feed hidden state $\mathbf{h}_{t}$ to an RMS normalization and a linear layer to compute a scalar value:
$g_{t}$ $=\mathtt{Linear}^{g}_{\phi}(\mathtt{RMSNorm}^{g}_{\phi}(\mathbf{h}_{t})),$ (23)
and the final output $\mathbf{o}_{t}$ is an interpolation between the FwPKM output and the value residual:
$\mathbf{o}_{t}$ $=g_{t}\cdot\mathtt{PKM}(\mathbf{q}_{t};\theta)+(1-g_{t})\cdot\mathbf{v}_{t}$ (24)
$=g_{t}\cdot\hat{\mathbf{v}}_{t+1}+(1-g_{t})\cdot\mathbf{v}_{t}.$ (25)
3.5 FwPKM Summary
We put everything together and show the complete formulation of FwPKM in the following and in Figure 1.
Figure 1: Architecture of FwPKM.
For each token $t$ in a chunk, slow-weight network computes inputs to FwPKM.
$\mathbf{q}_{t},\mathbf{v}_{t},g_{t}$ $=\mathtt{Linear}^{q,v,g}_{\phi}(\mathtt{RMSNorm}^{q,v,g}_{\phi}(\mathbf{h}_{t}))$ (26)
FwPKM’s forward pass predicts the value at $t+1$ from query at $t$, which is $\hat{\mathbf{v}}_{t+1}=\mathtt{FW\-PKM}(\mathbf{q}_{t};K^{\{1,2\}},V)$. A step-by-step process is as follows. We omit timestep subscript $t$ for variables other than the input $\mathbf{q}_{t}$ and output $\hat{\mathbf{v}}_{t+1}$.
$[\mathbf{q}^{1};\mathbf{q}^{2}]$ $=\mathbf{q_{t}}$ (27)
$s^{\{1,2\}}_{i}$ $=-\log(\epsilon+\|\mathbf{q}^{\{1,2\}}-K^{\{1,2\}}_{i}\|_{2}^{2})$ (28)
$\mathcal{I}^{\{1,2\}}$ $=\mathcal{T}_{k}(\mathbf{s}^{\{1,2\}})$ (29)
$\{s^{\{1,2\}\prime}_{i}\}$ $=\mathtt{softmax}(\{s^{\{1,2\}}_{j}\}_{j\in\mathcal{I}})$ (30)
$\mathcal{I}$ $=\mathcal{T}_{k}(\{s_{i}^{1}+s_{j}^{2}|i\in\mathcal{I}^{1},j\in\mathcal{I}^{2}\})$ (31)
$\hat{\mathbf{v}}_{t+1}$ $=\textstyle\sum_{i\in\mathcal{I}}V_{i}s^{\prime}_{i}$ (32)
The predicted value is combined with the value residual using the above gating value as weight. Lastly, an RMS normalization layer and a linear layer of slow weights are used to transform the output.
$\mathbf{o}_{t}$ $=g_{t}\cdot\hat{\mathbf{v}}_{t+1}+(1-g_{t})\cdot\mathbf{v}_{t}$ (33)
$\mathbf{o}^{\prime}_{t}$ $=\mathtt{Linear}^{o}_{\phi}(\mathtt{RMSNorm}^{o}_{\phi}(\mathbf{o}_{t}))$ (34)
Once we have collected the predicted values and target values for a chunk, we apply an update to the FwPKM’s parameters $\theta=\{V,K^{1},K^{2}\}$. The value matrix is updated using shaped gradients from the MSE losses weighted by gating values.
$\nabla^{\text{agg}}_{V_{i}}$ $=\frac{1}{N^{\text{read}}_{i}}\textstyle\sum_{t=1}^{C}\nabla_{V_{i}}\frac{1}{2}g_{t}\mathcal{L}_{\text{MSE}}(\hat{\mathbf{v}}_{t},\mathbf{v}_{t})$ (35)
$V^{\prime}_{i}$ $=V_{i}-\nabla^{\text{agg}}_{V_{i}}.$ (36)
The key matrices are updated using gradients from the marginal entropy-based addressing loss.
$\bar{\mathbf{p}}$ $=\frac{1}{C}\textstyle\sum_{t=1}^{C}\mathbf{s}^{\prime}_{t}$ (37)
$\mathcal{L}_{\text{addr}}$ $=\textstyle-\sum_{i=1}^{\sqrt{N}}\bar{p}_{i}\log\bar{p}_{i}$ (38)
$K^{\prime}$ $=K-\nabla_{K}\mathcal{L}_{\text{addr}},\quad K\in\{K^{1},K^{2}\}$ (39)
@@ 4.1 Training Setting
We implement language models based on the QwenNext architecture [^2] https://huggingface.co/Qwen/Qwen3-Next-80B-A3B-Thinking, which interleaves Gated DeltaNet (GDN, GDN) and gated softmax attention layers (GatedAttn). Our investigation uses these 12-layer baseline configurations.
@@
We introduce PKM and FwPKM modules into these baselines at specific depths (layers 2, 6, and 10). A standard PKM layer replaces the original FFN, while an FwPKM layer is inserted between the token mixer (Attention/GDN) and the FFN. Both PKM and FwPKM have $512^{2}$ memory slots. PKM retrieves effectively Top-$128$ slots ($4$ heads and $32$ slots per head) while FwPKM retrieves Top-$8$ slots ($1$ head and $8$ slots per head).
@@
We refer readers to Appendix A for more details of the architectures and optimization.

### v2

@@ 3 Fast-weight Product Key Memory
Figure 1: Architecture of FwPKM. An FwPKM layer is placed before a token mixer (i.e. a self-attention or a Gated DeltaNet layer). Slow-weight parameters are updated at only training time, using a global language modeling loss. Fast-weight value parameters and key parameters are updated at both training and inference time, by minimizing local memorization loss and addressing loss, respectively.
Standard Product Key Memory (PKM) is a slow-weight sparse channel mixer, meaning that its parameters are learned by the global language modeling objective over a corpus and then remain frozen at inference. As a result, PKM mainly stores semantic knowledge but cannot rapidly incorporate episodic information from the current input stream. We propose Fast-weight Product Key Memory (FwPKM), which keeps PKM’s retrieval process unchanged, but redesigns the memorization process by updating the memory matrices online. Concretely, FwPKM stores key-value associations from the input sequence by minimizing a local reconstruction objective over chunks, in the spirit of test-time training.
3.1 Overview
Figure 1 summarizes FwPKM. Given hidden states $\mathbf{h}_{1:T}$ [^1] We abuse tensor subscripts to index both sequence and feature dimensions; $t$ indexes sequence position and $i,j$ index feature dimensions., we process them in chunks of size $C$. Within each chunk, FwPKM: (i) constructs query/value vectors and a scalar gate using slow-weight projections, (ii) retrieves a sparse value prediction with PKM using the current fast weights, and (iii) writes to the fast weights by optimizing local objectives over the chunk. Concretely, we update the value matrix $V$ by minimizing a gated reconstruction loss $\mathcal{L}_{\text{mem}}$ (Section 3.3), and update the key matrices $K^{1},K^{2}$ by minimizing an auxiliary addressing loss $\mathcal{L}_{\text{addr}}$ that prevents slot collapse (Section 3.4). We apply these fast-weight updates after processing the chunk, so predictions within the chunk depend only on the fast weights from previous chunks, preserving causality.
3.2 Forward Pass: Inputs and Retrieval
Input construction.
For each token $t$, a slow-weight network parameterized by $\phi$ constructs inputs to FwPKM:
$\mathbf{q}_{t}$ $=\mathtt{Linear}^{q}_{\phi}(\mathtt{RMSNorm}^{q}_{\phi}(\mathbf{h}_{t})),$ (8)
$\mathbf{v}_{t}$ $=\mathtt{Linear}^{v}_{\phi}(\mathtt{RMSNorm}^{v}_{\phi}(\mathbf{h}_{t})),$ (9)
$g_{t}$ $=\sigma\!\left(\mathtt{Linear}^{g}_{\phi}(\mathtt{RMSNorm}^{g}_{\phi}(\mathbf{h}_{t}))\right).$ (10)
where $\mathbf{q}_{t}\in\mathbb{R}^{d_{k}},\mathbf{v}_{t}\in\mathbb{R}^{d_{v}},g_{t}\in(0,1)$ and $\sigma(\cdot)$ denotes the sigmoid function.
Sparse retrieval.
FwPKM uses the same product-key sparse retrieval as PKM (Section 2.2). Given $\mathbf{q}_{t}$, PKM returns a sparse index set $\mathcal{I}_{t}$ and normalized weights $\{s^{\prime}_{t,i}\}_{i\in\mathcal{I}_{t}}$, and predicts a value by a weighted sum of value rows:
$\hat{\mathbf{v}}_{t}=\mathtt{PKM}(\mathbf{q}_{t};\theta)=\sum_{i\in\mathcal{I}_{t}}s^{\prime}_{t,i}\,V_{i}.$ (11)
Gated Residual Output.
Next-token prediction does not always require episodic memory, so we interpolate the retrieved value with a residual value path:
$\mathbf{o}_{t}=g_{t}\cdot\hat{\mathbf{v}}_{t}+(1-g_{t})\cdot\mathbf{v}_{t},\qquad\mathbf{o}^{\prime}_{t}=\mathtt{Linear}^{o}_{\phi}(\mathtt{RMSNorm}^{o}_{\phi}(\mathbf{o}_{t})).$ (12)
3.3 Backward Pass: Memorization by Learning to Reconstruct
Chunk-level local objective.
Within each chunk, FwPKM updates fast weights $\theta$ to make retrieval reconstruct target values derived from the same stream. Following TTT-style fast weights, we minimize a reconstruction loss over query-value pairs. For a chunk of size $C$, define
$\mathcal{L}_{\text{mem}}$ $=\sum_{t=1}^{C}\frac{1}{2}\,g_{t}\,\|\mathbf{v}_{t}-\hat{\mathbf{v}}_{t}\|_{2}^{2}.$ (13)
Why MSE and the $\tfrac{1}{2}$ factor.
[…]
Gradient aggregation.
Multiple tokens in a chunk may write to the same value row. Let $N_{i}^{\text{read}}$ denote the number of times row $V_{i}$ is accessed in the chunk. We aggregate per-row gradients by normalizing with $N_{i}^{\text{read}}$. This aggregation works as a consensus mechanism for competing memory writings.
$\nabla^{\text{agg}}_{V_{i}}=\frac{1}{N^{\text{read}}_{i}}\nabla_{V_{i}}\mathcal{L}_{\text{mem}}.$ (14)
For the memorization, we use $\nabla^{\text{agg}}_{V_{i}}$ to update the accessed value rows:
$V_{i}\leftarrow V_{i}-\nabla^{\text{agg}}_{V_{i}}.$ (15)
3.4 Backward Pass: Prevent Slot Collapsing via Addressing Optimization
Sparse memories can suffer from memory slot collapsing, where only a small fraction of slots are used. We therefore optimize an auxiliary addressing objective that encourages balanced slot usage on average across a chunk, without forcing each individual query to be uniform.
Let $\mathbf{s}^{\prime 1}_{t}\in\mathbb{R}^{\sqrt{N}}$ and $\mathbf{s}^{\prime 2}_{t}\in\mathbb{R}^{\sqrt{N}}$ be the normalized Top-$k$ score vectors over the two sub-key sets (unselected indices have score 0). Define marginal slot-usage distributions as:
$\bar{\mathbf{p}}^{1}=\frac{1}{C}\sum_{t=1}^{C}\mathbf{s}^{\prime 1}_{t},\qquad\bar{\mathbf{p}}^{2}=\frac{1}{C}\sum_{t=1}^{C}\mathbf{s}^{\prime 2}_{t},$ (16)
and minimize $\mathcal{L}_{\text{addr}}$ which is the negative entropy:
$\mathcal{L}_{\text{addr}}$ $=-H(\bar{\mathbf{p}}^{1})-H(\bar{\mathbf{p}}^{2})$
$=-\sum_{i=1}^{\sqrt{N}}\bar{p}^{1}_{i}\log\bar{p}^{1}_{i}\;-\;\sum_{i=1}^{\sqrt{N}}\bar{p}^{2}_{i}\log\bar{p}^{2}_{i}.$ (17)
We update keys using the gradients of $\mathcal{L}_{\text{addr}}$:
$K^{1}\leftarrow K^{1}-\nabla_{K^{1}}\mathcal{L}_{\text{addr}},\qquad K^{2}\leftarrow K^{2}-\nabla_{K^{2}}\mathcal{L}_{\text{addr}}.$ (18)
3.5 Practical Choices
We found the following choices helpful for improving model performance and training stability. Detailed discussion and ablations are deferred to Appendix A and Appendix D, respectively.
  • Lookahead value targets. During memorization, we pair $\mathbf{q}_{t}$ with $\mathbf{v}_{t+1}$. In this way, a query extracts information about the immediate future of similar queries in history, which aligns the local reconstruction task with the global next-token prediction task (Appendix A.1).
  • IDW scoring. Instead of scoring a key row by dot product $\mathbf{q}^{(m)\top}K^{(m)}_{i},m\in\{1,2\}$ (Eq. 5), we use an inverse-distance weighting (IDW, IDW) score that depends on Euclidean distance, which encourages keys to become clustering centroids in representation space. We apply IDW by changing the scoring function from $\mathbf{q}^{\top}K_{i}$ to $\log\big(\epsilon+\|\mathbf{q}-K_{i}\|_{2}^{2}\big)$ (Appendix A.2).
  • Target normalization. We z-score normalize $\mathbf{v}_{t}$ along the feature dimension for stability (Appendix A.3).
@@ 4 Experiments
We evaluate FwPKM along four axes. First, we use perplexity on short- and long-context corpora to disentangle semantic (slow-weight) and episodic (fast-weight) memory roles. Next, we stress-test episodic retrieval with Needle-in-a-Haystack, including iterative rereading that leverages test-time optimization. Finally, we evaluate on realistic long-context QA tasks from LongBench and probe online domain adaptation on diverse Pile subsets.
@@ 4.1 Training Setting
We implement 12-layer language models that follow the QwenNext design principles (qwen3technicalreport; Qwen3Next80BThinking) but are scaled down to our setting (baseline models are $\sim$112M parameters), interleaving Gated DeltaNet (GDN, GDN) and gated softmax attention layers (GatedAttn). Our investigation uses the following baseline configurations.
@@
We introduce PKM and FwPKM modules into these baselines at specific depths (layers 2, 6, and 10). A standard PKM layer replaces the original MLP, while an FwPKM layer is inserted before the token mixer (Attention/GDN). Both PKM and FwPKM have $512^{2}$ memory slots (262,144262{,}144). PKM retrieves effectively Top-$128$ slots ($4$ heads and $32$ slots per head) while FwPKM retrieves Top-$8$ slots ($1$ head and $8$ slots per head). We choose $8$ slots because it hits a balance between efficiency and performance. See ablation experiments in Appendix D.
@@
We refer readers to Appendix C for more details of the architectures and optimization.